Will someone please help me with this story problem that has been a struggle for me.
Trains A and B are traveling in the same direction on parrellel tracks. Train A is traveling at 100 miles per hour and train B is traveling at 125 miles per hour. Train A passes a station at 2:25PM. If train B passes the same station at 2:37PM at what time will train B catch up to train A?
When will train B catch up to train A? ___:___ am or pm?
Thanks for your help.

Answer by mananth(16949)

Trains A and B are traveling in the same direction on parrellel tracks. Train A is traveling at 100 miles per hour and train B is traveling at 125 miles per hour. Train A passes a station at 2:25PM. If train B passes the same station at 2:37PM at what time will train B catch up to train A?
When will train B catch up to train A? ___:___ am or pm?
difference in time when the trains passes the station is 12 minutes = 1/5 hours.
speed of train B = 125 mph
speed of Train A = 100 mph
In 1/5 hour train A has traveled 1/5 *100 = 20 miles.
difference in speed = 25 mph.
Train B will catch up 25 miles every hour.
To catch up 20 miles it will take 20/25 hours
=4/5 hours = 48 minutes
So Train B will catch up with train A at 2.37 + 0.48 minutes = 3.25 pm